OpenStudy (jazzyluvsu123):

a recording song can last 2.5 minutes. What is the Greatest number of times the song fully be played in a hour ? best reponse get medal ! @VoxHavoc @Directrix

Directrix (directrix):

@JazzyLuvsU123 An hour is 60 minutes so I'm thinking that if you divide 60 by 2.5 then the answer will be the number of times the record can be played. wHAT IS 60/2.5 =

OpenStudy (jazzyluvsu123):

ok I will try will quick

OpenStudy (anonymous):

You would have to divide 60 by 2.5 so \[\frac{ 60 }{ 2.5 }=?\]

Directrix (directrix):

@JazzyLuvsU123 Put this "60/2.5 =" in the Google search box and the Google calculator will pop up.

OpenStudy (jazzyluvsu123):



Directrix (directrix):

Nope. Do this: @JazzyLuvsU123 Put this "60/2.5 =" in the Google search box and the Google calculator will pop up with the answer.

OpenStudy (anonymous):

You have to do it the other way.